OpenStudy (anonymous):

A zoo charges admission of $4.75 for adults and $2.75 for children. A group of 12 people paid for admission with a $50- dollar bill. If the group received at least $9.00 in change, what is the maximum number of adults in the group ? PLEASE HELP !!!

OpenStudy (anonymous):

Are there any answer choices?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

hmm okay. do you have anything so far?

OpenStudy (anonymous):

Uh- no

OpenStudy (anonymous):

Okay - I'll try to help and explain.

OpenStudy (anonymous):

I didn't really understand the question that well.

OpenStudy (anonymous):

Okay so these 12 people give a $50 bill and got $9.00 back in change. So that means the price of all their tickets would have to be $41.00, right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

this can be represented as 2.75x+4.75y=41 x+y=12 Where x is the children and y is the adults

OpenStudy (anonymous):

^yep

OpenStudy (anonymous):

This is a system of equations and can be solved as such How could you solve this?

OpenStudy (anonymous):

I'm not sure but how about subtracting 2.75 from 4.75 which would give me 2.00 then divide it by 41 @Brill

OpenStudy (anonymous):

Well... you can't subtract them because they are not like terms... Are you allowed the use of a calculator?

OpenStudy (anonymous):

what Mathematics class is this for? are you allowed to use calc

OpenStudy (anonymous):

A calculator would make this easier, more specifically a graphing calculator

OpenStudy (welshfella):

youy can use the substitution method to solve this from the 2nd eqaution y = 12-x Plug this in for y in first equation and solve for x.

OpenStudy (anonymous):

But if are not allowed to use a calculator you can solve this through substitution, By manipulating x+y=12 You can see that y=12-x So substitute 12-x for y in the first equation getting 2.75x+4.75(12-x)=41

OpenStudy (welshfella):

then you can find y which will equal 12 - x

OpenStudy (anonymous):

\[2.75x+57-4.75x=41\]

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

can you continue?

OpenStudy (welshfella):

add like terms on left side

OpenStudy (anonymous):

Wait- do I combine -4.75 and 2.75

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

I did that and i got 7.5

OpenStudy (anonymous):

So now I have 7.5x+57=41

OpenStudy (welshfella):

no the answer must be negative because 4.75 is greater than 2.75

OpenStudy (welshfella):